OpenStudy (anonymous):

5x-2=28-x I need to see the steps for this problem please. I already know the answer is 5. Please and thank you.

OpenStudy (anonymous):

5x-2=28-x put the x into 1 side, how are you going to do that?

OpenStudy (anonymous):

Add the x to both sides?

OpenStudy (anonymous):

aha, do it

OpenStudy (anonymous):

5x-2=28

OpenStudy (anonymous):

u said add x to both sides, however u added it to only one

OpenStudy (anonymous):

Ah, now I'm just confused.

OpenStudy (gudden):

@AlexusFerilli um... i have another method of solving it... can i try explaining it ??

OpenStudy (anonymous):

@gudden Go ahead.

OpenStudy (anonymous):

5x-2=28-x add x to both sides 5x-2 + x=28-x +x 6x - 2 = 28 now move the 2; u add 2 to both sides 6x - 2 + 2 = 28 +2 6x = 30 divide both by 6 6x/6 = 30 /6 x = 5

OpenStudy (anonymous):

Thank you! ♥♥♥

OpenStudy (gudden):

@AlexusFerilli u can also take the x to a single side and the numbers to the other

OpenStudy (gudden):

the question given is 5x - 2 = 28 - x on taking x to a single side and the numbers to the other we get 6x = 30 now x = 30/6 = 5